OpenStudy (anonymous):

Two numbers total 6, and their difference is 4. Find the two numbers. PLEASE EXPLAIN!! :(

jimthompson5910 (jim_thompson5910):

Let x and y be your two numbers. They are unknown for now.

jimthompson5910 (jim_thompson5910):

"Two numbers total 6" means that they add to 6. So, x+y = 6 "their difference is 4" means the result after subtraction is 4. That equation is x-y = 4

jimthompson5910 (jim_thompson5910):

So you have this system of equations x+y = 6 x-y = 4 We can use elimination to solve. Add the equations together to get x+y = 6 x-y = 4 ------- 2x+0y = 10 Since 2x = 10, we know that x = ______

OpenStudy (anonymous):

5? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

then you use x to find y

jimthompson5910 (jim_thompson5910):

plug x = 5 into any equation with x,y in it. Then solve for y

OpenStudy (anonymous):

thank you very much!! @jim_thompson5910 :D

jimthompson5910 (jim_thompson5910):

np